OpenStudy (anonymous):

if x^2=27 what is the value of x^4/3?

OpenStudy (fibonaccichick666):

so first, do you know what x^(4/3) means?

OpenStudy (anonymous):

Well, first find the value of x, and then substitute it in for the x in the next problem!

OpenStudy (fibonaccichick666):

(I'm assuming no calculators are allowed, though both approaches work)

OpenStudy (gorv):

x^4=(x^2)^2

OpenStudy (gorv):

orx^4=x^2*x^2

OpenStudy (gorv):

i think u can proceed now??

OpenStudy (gorv):

@dawnum08

OpenStudy (anonymous):

i don't understand...how does that relate to x^2=27?

OpenStudy (anonymous):

well, you know it is basically saying x times itself equals 27.. think which number times itself equals 27

OpenStudy (fibonaccichick666):

can you explain what x^(4/3) means irst, then I promise I'll explain how it relates

OpenStudy (anonymous):

i think i have it, x^2=27, take the square root of both sides you get x=+-3(squareroot of 3), plug both of those in for x in x^4/3?

OpenStudy (fibonaccichick666):

no, that would be the cubed root

OpenStudy (anonymous):

how do you write a square root symbol on a keyboard?

OpenStudy (fibonaccichick666):

please, first, what does x^(4/3) mean

OpenStudy (fibonaccichick666):

I apologize, I missed your parentheses, thst is correct but much harder than necessary

OpenStudy (fibonaccichick666):

so please, first, tell me what x^(4/3) means

OpenStudy (anonymous):

i don't understand what you're asking? it means just what it says x to the 4/3 power.

OpenStudy (anonymous):

4/3 is the exponent.

OpenStudy (anonymous):

hello?

OpenStudy (fibonaccichick666):

right, but let's say you had 2^(4/3) how would you compute that?

OpenStudy (fibonaccichick666):

without putting it into a calculator

OpenStudy (fibonaccichick666):

ie. what does a rational exponent mean?

OpenStudy (anonymous):

i don't know.

OpenStudy (fibonaccichick666):

please read this for me: http://www.purplemath.com/modules/exponent5.htm

OpenStudy (fibonaccichick666):

once you know about them, it should click as to what @gorv was hinting at with his/her information

OpenStudy (anonymous):

i'm confused, is my answer wrong?

OpenStudy (fibonaccichick666):

Yes.

OpenStudy (gorv):

\[x^{\frac{ 4 }{ 3 }}=(x^{4})^{\frac{ 1 }{ 3 }}=(x^2*x^2)^{\frac{ 1 }{ 3 }}\]

OpenStudy (fibonaccichick666):

Now, did you read the website?

OpenStudy (fibonaccichick666):

if you did, what gorv did ^^ should make sense

OpenStudy (anonymous):

x^4/3 = the cube root of x^4, i get that

OpenStudy (fibonaccichick666):

ok, so, do you understand that\( x^4 =x^2x^2\)

OpenStudy (fibonaccichick666):

and can you tell me why?

OpenStudy (anonymous):

yes, i understand (27*27)^1/3

OpenStudy (fibonaccichick666):

ok, so, can you simplify that? (1 number)

OpenStudy (fibonaccichick666):

then you are done. It wanted the value of \(x^{4/3}\) not the value of x.

OpenStudy (anonymous):

thank you.

OpenStudy (fibonaccichick666):

np